OpenStudy (anonymous):

Find the 9th term of the arithmetic sequence. a3 = 12.2 and a4 = 12.6; a9 = _____.

OpenStudy (experimentx):

d = a4 - a3 = 0.4 a3 = a1+2d => a1 = 12.2 - 0.4*2 = 11.4 a9 = 11.4+8*0.4